OpenStudy (anonymous):

what is its dy/dx without simplifying the answer? 1) sin x + sin y = 1 step by step please

OpenStudy (cwrw238):

implicit differentiation you treat y as a function of x cos x + cos y * dy/dx = 0 now make dy/dx the subject

OpenStudy (cwrw238):

I've applied the chain rule to sin y

OpenStudy (anonymous):

sorry but still confused

OpenStudy (cwrw238):

Ok - you find the derivative term by term derivative wrt x for sin x = cos x derivative for the constant 1 is 0 ok?

OpenStudy (anonymous):

wrt?

OpenStudy (cwrw238):

with respect to

OpenStudy (anonymous):

oh ok

OpenStudy (cwrw238):

now for sin y we treat y as a function of x. The equation implies that.

OpenStudy (anonymous):

so d/dx (sin(y(x))) = d/dx (1) then we differentiate term by term?

OpenStudy (cwrw238):

suppose y = x^2 we'd have sin x^2 and the derivative of that , by the chain rule is d(x^2) / dx * d(sin x^2) / dx = 2x * cos x^2

OpenStudy (cwrw238):

but we have y not x^2 so we write dy/dx and cos y

OpenStudy (anonymous):

ok thanks i got it

OpenStudy (cwrw238):

yw